OpenStudy (anonymous):

Find the area of the sector if m∠ A = π/10 and the circle has a radius of 20 inches.

OpenStudy (anonymous):

Do you mean \[\frac{ \pi }{ 10 }\]?

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Well, this is a sector, so it has to be a part of a sircle with radius 20 inches. It's obvious that the radius will be the same, but the circumference/perimeter of the sector and the entire sircle will differ. The relation between the circumferences is\[\frac{ \frac{ \pi }{ 10 } }{ 2\pi }\]\[= 1/20\]. Therefore, the area of the sector must as well be 1/20 of the area of the entire sircle. A sircle with radius of 20 inches has area = 20*20*pi = 400 pi. Then the sector with angle pi/10 has to have area 400pi/20 = 20pi.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

mind helping with another problem ?

OpenStudy (anonymous):

Ok, bring it on

OpenStudy (anonymous):

Calculate the average rate of change of the function f(x) = mx + b over the interval [2, 13].

OpenStudy (anonymous):

Have you learned about the derivative?

OpenStudy (anonymous):

no i have not

OpenStudy (anonymous):

Ok, then I will try to write it as easy as I can. The average rate of change of the function on the interval [2, 13] has to be the average change in y, or f(x) over the interval of x. As you can see, the difference in x = 13-2 = 11. What about the change in y, or f(x)? Well it has to be f(13) - f(2) = m*13 + b - (m*2 + b) = 13*m + b - 2*m - b = 11*m. Then, the average rate of change of the x interval [2, 13] is often written as change in y divided by change in x, or \[\frac{ \Delta y }{ \Delta x } = \frac{ \Delta f(x) }{ \Delta x } = \frac{ 11m }{ 11 } = m\]

OpenStudy (anonymous):

ahh i understand , you are very helpful thank you once again

OpenStudy (anonymous):

No problem, glad I could help :)